Two cards are drawn from a standard deck of cards. What is the probability that:
a. both are jacks and ace cards?
b. both are diamonds?
c. both are spades or face cards?
d. one is black and one is red?
e. at least one card is red?

Answer by Alan3354(69443)

Two cards are drawn from a standard deck of cards. What is the probability that:
----
If you mean without replacement:
a. both are jacks and ace cards?
Confusing wording. If you mean 1 J and 1 Ace:
The 1st draw is 4/52 = 1/13
The 2nd draw is 4/51
(1/13)*(4/51) = 4/663
-------
The other problems are similar.
Try them, email via the TY note for help or to check your work.
=========================================
b. both are diamonds?
c. both are spades or face cards?
d. one is black and one is red?
e. at least one card is red?